# Ab Initio Framework of Electron-Phonon Coupling for Chiral Phonons with Giant Phonon Magnetic Moments in Magnetic Materials

### v1

@@ Abstract
Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention. However, this field remains severely limited due to the inability to accurately account for electron-phonon-coupling (EPC) effects for chiral phonons. Here, we resolve this challenge through developing an ab initio framework for computing EPC-driven phonon magnetic properties, applicable to all insulating and metallic materials. Our theory predicts a dramatic enhancement of phonon magnetic moments exceeding classical predictions by multiple orders, accompanied by distinct Zeeman splitting. Interestingly, these splittings induce observable topological gaps in phonon spectra of magnets, generating intrinsic phonon Chern states. Through first-principles calculations, we identify five topological phonon magnets with intrinsic phonon Chern states and also propose to detect low-energy neutral particles such as dark matters. This research not only establishes a computational framework for investigating EPC-driven phonon magnetic properties, but also provides a material design strategy for realizing topological phonon states in quantum magnetic materials.
@@
Introduction. Chiral phonons, defined as collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and his cooperators [1, 2] and subsequently verified experimentally in monolayer WSe2 [3]. These chiral vibrations not only reveal fundamental properties of phonons previously unrecognized, but also establish a novel degree of freedom for manipulating phonon-mediated phenomena [1, 4, 5, 6, 7]. Recently, there has been a surge of research interest in chiral phonons, spanning theoretical investigations [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] to experimental explorations [19, 20, 21, 22, 23, 24, 25, 26, 27]. Of particular importance are phonon magnetic moments (PMMs) and their associated phonon Zeeman splittings, which constitute key mechanisms in chiral-phonon-involved effects [6, 7]. However, classical models based on point-charge approximations yield PMM values that are orders of magnitude below practical application requirements [13], posing significant limitations for the advancing development and application of chiral phonons. Although electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [11, 13, 16], a unified ab initio theoretical framework of EPC-driven PMMs for realistic materials is lacking.
@@
Figure 1: Schematic illustrations of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-driven PMM. In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions, which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons. The classical Zeeman splitting of chiral phonons is negligible as a result of the significant atomic mass, whereas EPC-deiven Zeeman splitting is orders of magnitude larger. This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
@@
Besides magnetic moments, phonons can also demonstrate the Hall effect [28, 29, 30, 31] and manifest topological characteristics, such as Berry curvature and Chern numbers [32, 33, 34]. When the time reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, essentially realizing phonon Chern insulators. These topological phonon states featuring topologically protected edge currents could significantly enhance both detection sensitivity and practical exploitation of the phonon Hall effect. Nevertheless, conventional TRS-breaking approaches for phonons, such as strong external magnetic fields or Coriolis forces, often face experimental limitations or induce parasitic effects [35, 36, 37]. Consequently, phonon Chern states driven by intrinsic magnetic moments without external fields are highly desirable. Crucially, predicting such intrinsic topological phonon states in magnets demands precise calculations of TRS-broken phonon spectra, which remains challenging due to the lack of ab initio framework for magnetic phonon systems [38, 39, 40, 41, 42, 43].
@@
To resolve these challenges, we establish a unified ab initio framework for determining EPC-driven phonon magnetic properties, which enables first-principles calculations of the previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals. Our theory demonstrates EPC-driven enhancements of PMMs by four orders of magnitude, inducing giant chiral phonon Zeeman splittings. Building on these giant splittings, we propose the inertially decoupled model to realize intrinsic phonon Chern states in magnets. Furthermore, through first-principles calculations, we identify five promising candidate materials exhibiting intrinsic phonon Chern states, including $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ and $\mathrm{CrB_{2}}$. Our work not only paves a material-design pathway for topological phonon magnets, but also establishes the framework for chiral phonon physics in magnetic phenomena.
@@
Ab initio framework for phonon Zeeman energy. Previous theories regarding the magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals and are only effective for specific high-symmetry phonon wave vectors [11, 13, 14, 16]. However, since EPC is typically much more pronounced in metallic systems, there is a pressing need to develop a unified theory that can comprehensively describe phonon behaviors in insulators, semimetals, and metals.
@@
We begin with a TRS-breaking perturbation phonon Hamiltonian, $\delta H=-\sum_{ab}\eta_{ab}p_{a}u_{b}$, known as the Raman spin-phonon interaction [44], where $\eta_{ab}$ is a real antisymmetric matrix that satisfies $\eta_{ab}=-\eta_{ba}$. Here, $u_{a}=\sqrt{m}x_{a}$ represents the renormalized atomic displacement coordinate and $p_{a}$ represents the canonical momentum related to $u_{a}$, where $m$ denotes the atomic mass. By identifying the phonon angular momentum as $s^{\text{ph}}_{c}=\sum_{ab}\epsilon_{abc}u_{a}p_{b}$ (c is the direction of the phonon angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express the phonon Zeeman energy in a compact form as $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$. Here, we have defined $\eta_{c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{ab}$. Therefore, the Zeeman energy of the chiral phonon can be regarded as an effective linear response to its angular momentum.
@@
We notice that $\bm{\eta}$ can be deduced from the TRS-breaking self-energy correction of chiral phonons driven by EPC. Remarkably, within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where the $\chi_{c}^{-}(\bm{q},\omega)$ component is defined as $\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$. The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig. 1(b)(see the supplement material (SM)) [45],
@@
Large phonon Zeeman splittings suggest that phonons exhibit significant phonon magnetic moments. To derive the PMM, we taking the derivative $\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}$ and the phonon Zeeman energy is further rewritten as $E_{\text{Z}}=-\bm{B}\cdot\bm{\mu}^{\text{ph}}$, where $\bm{B}$ represents the effective magnetic field, and $\bm{\mu}^{\text{ph}}$ denotes the PMM (see SM  [45]),
@@
Candidate materials. Building upon the above theoretical framework, we developed a dedicated computational module within Quantum ESPRESSO [52, 53] incorporating density functional perturbation theory with Hubbard corrections (DFPT+U) [46, 47] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons. Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from dynamical matrix through near-degenerate perturbation theory (see SM [45]), yielding TRS-broken phonon dispersions.
@@
Discussion and conclusion. The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring extreme external strong magnetic fields, Coriolis forces, or stress gradients. One potential advantage may be its ability to detect neutral particles by converting them into phonon edge currents. Neutral particles, especially dark matter particles, rarely participate in electromagnetic interactions, making their detection extremely challenging. However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within that of the first “conduction” band above the topological gap in the phonon dispersion, this process may potentially excite the corresponding phonon modes [64, 65], driving the system out of equilibrium. As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via edge states, producing a edge phonon current immune to scattering[Fig. 4(a)]. This robust edge current is significantly more measurable than conventional phonon excitations dispersed throughout the material. However, the operating temperature of detectors must be sufficiently low to prevent thermal excitation from overcoming the phonon gap, which would otherwise disrupt these edge currents. Thus, the gap determines the maximum operating temperature for such a detector. Fig. 4(b) lists the transfered energy ranges and gaps for all five candidate materials (see SM [45] for details), with transfered energy spanning approximately $20\sim 110\,\mathrm{meV}$. This energy range corresponds to light-mass dark matter, therefore our work provides a potential detection mechanism for this recently emphasized category of dark matter candidates.
@@
In summary, we have established a unified ab initio framework enabling systematic predictions of EPC-driven phonon magnetic properties and TRS-breaking phonon dispersions across for insulator, semimetal and metal systems. Under this framework, we propose an inertially decoupled model to design intrinsic phonon Chern states, and identify five candidate materials exhibiting EPC-driven giant topological phonon gaps exceeding 0.3 THz. Furthermore, we demonstrate the potential capability of these topological chiral phonon states to detect neutral particles through phonon edge current transduction. These advances simultaneously open broad pathways for revealing new phonon magnetic effects in quantum magnetic materials.

### v2

# Ab Initio theory of Electron-phonon-coupling-induced Giant Magnetic Moments of Chiral Phonons in Magnetic Materials

@@ Abstract
Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention. However, a long-standing critical issue in this field is the lack of ab initio method to accurately calculate phonon magnetic moments resulting from electron-phonon coupling (EPC). Here, we resolve this challenge by developing an ab initio theory for calculating EPC-induced phonon magnetic properties, applicable to both insulating and metallic materials. Based on this theory, we demonstrate EPC-induced giant phonon magnetic moments and resulting phonon Zeeman splittings in magnetic metals, which are orders of magnitude larger than classical predictions from point-charge models. Interestingly, these splittings could open observable topological gaps in phonon spectra of magnets, generating intrinsic phonon Chern states. Through first-principles calculations, we first propose candidate materials with such intrinsic phonon Chern states hosting robust edge phonon currents which may be applied to detecting neutral particles such as dark matter particles. Our work not only establishes a theoretical foundation for EPC-induced phonon magnetic properties, but also enables the ab initio calculation of long-sought TRS-breaking phonon spectra throughout the Brillouin zone in realistic materials.
@@
Introduction. Chiral phonons, collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and Niu [1, 2] and subsequently verified experimentally in monolayer WSe2 [3]. Recently, there has been a surge of research interest in chiral phonons, covering both theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and experiments [15, 16, 17, 18, 19, 20, 21, 22, 23]. Of particular importance are phonon magnetic moments (PMMs) of chiral phonons and their associated phonon Zeeman splittings, playing a crucial role in chiral-phonon-involved effects [1, 24, 25, 26, 27]. However, classical models based on point-charge approximations yield extremely small PMM values due to the significant large mass of nucleus  [9]. While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [7, 9, 12], an ab initio method capable of accurately calculating EPC-induced PMMs in general materials is still lacking.
@@
Figure 1: Schematic illustrations of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-induced PMM. In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions, which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons. The classical Zeeman splitting of chiral phonons deduced from the point-charge model is negligible as a result of the significant atomic mass, whereas EPC-deiven Zeeman splitting is orders of magnitude larger. This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
@@
Besides magnetic moments, phonons can also exhibit Hall effects [28, 29, 30, 31] and topological characteristics, such as Berry curvature and Chern numbers [32, 33, 34]. When the time reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, featuring topologically protected phonon edge currents with significantly enhanced detectability compared to conventional phonon Hall effects. Nevertheless, TRS-breaking external fields for phonons, such as strong external magnetic fields or Coriolis forces, often face experimental limitations [35, 36]. Consequently, phonon Chern states driven by the phonon Zeeman splitting in magnets without external fields are highly desirable. However, predicting such intrinsic topological phonon states in magnets remains challenging due to the inability to perform ab initio calcualtions of TRS-broken phonon spectra [37, 38, 39, 40, 41, 42].
@@
To resolve these challenges, we establish an ab initio theory for determining EPC-induced phonon magnetic properties, which enables first-principles calculations of previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals. Our theory relies on treating the angular momentum of chiral phonon as an effective perturbative field to the electron-phonon coupled system, with the phonon Zeeman energy interpreted as the response to this perturbation. Based on this theory, we demonstrates EPC-induced giant PMMs leading to large phonon Zeeman splittings in magnetic metals, which are orders of magnitude stronger than classical predictions from point-charge models. Building on these splittings, we propose the inertially decoupled model to search for intrinsic phonon Chern states in realistic magnetic materials. Furthermore, through first-principles calculations, we propose promising candidate materials exhibiting intrinsic phonon Chern states, including $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ and $\mathrm{CrB_{2}}$. Our work establishes a ab initio theoretical foundation for investigating EPC-induced phonon magnetic properties and TRS-breaking topological phonon spectra in realistic materials, which could pave the way for futher developments of chiral phonon physics.
@@
Ab initio theory for phonon Zeeman energy. Previous theories regarding the magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals and are only effective for specific high-symmetry phonon wave vectors [7, 9, 10, 12]. However, since EPC is typically much more pronounced in metallic systems, there is a pressing need to develop an ab initio theory that can comprehensively describe phonon magnetic properties in insulators, semimetals, and metals.
@@
Instead of focusing on calculating effective PMM, we choose to directly calculate experimentally observable Zeeman splittings. We begin with a TRS-breaking perturbation phonon Hamiltonian, $\delta H=-\sum_{ab}\eta_{ab}p_{a}u_{b}$, known as the Raman spin-phonon interaction [43], where $\eta_{ab}$ is a real antisymmetric matrix that satisfies $\eta_{ab}=-\eta_{ba}$. Here, $u_{a}=\sqrt{m}x_{a}$ represents the renormalized atomic displacement coordinate and $p_{a}$ represents the canonical momentum related to $u_{a}$, where $m$ denotes the atomic mass. By identifying the phonon angular momentum as $s^{\text{ph}}_{c}=\sum_{ab}\epsilon_{abc}u_{a}p_{b}$ (c is the direction of the phonon angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express the phonon Zeeman energy in a compact form as $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$, where we have defined $\eta_{c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{ab}$. Based on the expression of phonon Zeeman energy, we propose that the angular momentum of chiral phonon can be treated as an effective perturbative field to the electron-phonon coupled system. Therefore, the Zeeman energy of chiral phonon can be regarded as an effective linear response to its angular momentum.
@@
Remarkably, $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of chiral phonons induced by EPC. Within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where the $\chi_{c}^{-}(\bm{q},\omega)$ component is defined as $\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$. The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig. 1(b)[see the supplement material (SM)] [44],
@@
Large phonon Zeeman splittings suggest that phonons exhibit significant PMM. To derive the PMM, we taking the derivative $\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}$ and the phonon Zeeman energy is further rewritten as $E_{\text{Z}}=-\bm{B}\cdot\bm{\mu}^{\text{ph}}$, where $\bm{B}$ represents the effective magnetic field, and $\bm{\mu}^{\text{ph}}$ denotes the PMM (see SM  [44]),
@@
Candidate materials. Building upon the our theory, we developed a dedicated computational module within Quantum ESPRESSO [51, 52] incorporating density functional perturbation theory with Hubbard corrections (DFPT+U) [45, 46] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons. Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from dynamical matrix through near-degenerate perturbation theory (see SM [44]), yielding TRS-broken phonon dispersions.
@@
Discussion and conclusion. The intrinsic phonon Chern state, which hosts topologically protected edge phonon currents, holds significant potential for applications due to its operation without requiring extreme external strong magnetic fields, Coriolis forces, or stress gradients. One potential advantage may be its ability to detect neutral particles by converting them into phonon edge currents. Neutral particles, especially dark matter particles, rarely participate in electromagnetic interactions, making their detection extremely challenging. However, if the energy transferred from the neutral particle to nucleus during the scattering or absorption falls within that of the first “conduction” band above the topological gap in the phonon dispersion, this process may potentially excite the corresponding phonon modes [63, 64], driving the system out of equilibrium. As the system relaxes back to equilibrium, these high-frequency modes gradually shift to lower frequencies and necessarily cross the phonon Chern gap via edge states, producing a edge phonon current immune to scattering[Fig. 4(a)]. This robust edge current is significantly more measurable than conventional phonon excitations dispersed throughout the material. However, the operating temperature of detectors must be sufficiently low to prevent thermal excitation from overcoming the phonon gap, which would otherwise disrupt these edge currents. Thus, the gap determines the maximum operating temperature for such a detector. Figure 4(b) lists gaps and transfered energy ranges for proposed candidate materials (see SM [44] for details), with transfered energy spanning approximately $20\sim 110\,\mathrm{meV}$. This energy range corresponds to low-energy dark matter, therefore our work provides a potential detection mechanism for this recently emphasized category of dark matter candidates [63, 64].
@@
In summary, we have established an ab initio theory enabling systematic predictions of EPC-driven phonon magnetic properties and TRS-breaking phonon dispersions across for insulator, semimetal and metal systems. Based on our theory, we build the inertially decoupled model to seek intrinsic phonon Chern states, and propose candidate materials exhibiting EPC-driven giant topological phonon gaps exceeding 0.3 THz. Furthermore, we demonstrate the potential capability of these topological chiral phonon states to detect neutral particles through robust phonon edge currents. These advances could open broad pathways for investigating phonon magnetic effects in quantum materials.